OpenStudy (anonymous):

find the limit as x approaches 0,( x-sinx)/x^3

OpenStudy (mathmath333):

apply the same l'hospital rule

OpenStudy (anonymous):

differentiating will give \[\frac{ 1-\cos x }{ 3x^{2} }\]

OpenStudy (mathmath333):

yes,apply the same l'hospital rule once again

OpenStudy (anonymous):

substituting x=0, \[\frac{ 1-\cos0 }{ 3(0)^{2} }=\frac{ 1-1 }{ 0 }=\frac{ 0 }{ 0}=0\]

OpenStudy (mathmath333):

\(\large \lim_{x\to\ 0}\frac{(x-sinx)}{x^3}\) \(\large =\dfrac{d}{dx}\left(\dfrac{x-sinx}{x^3}\right)\) \(\large =\dfrac{1-cosx}{3x^2}\) \(\large =\dfrac{d}{dx}\left(\dfrac{1-cosx}{3x^2}\right)\) \(\large =\dfrac{-(-sinx)}{6x^1}\) \(\large =\dfrac{d}{dx}\left(\dfrac{sinx}{6x^1}\right)\) \(\large =\dfrac{cosx}{6}\) \(\large =\dfrac{1}{6} (as~~ cos0=1)\)

OpenStudy (anonymous):

how many times do i have to differentiate?

OpenStudy (mathmath333):

sorry three times here

OpenStudy (anonymous):

how do i know that i have to end on the first,second ,third, forth etc. differential

OpenStudy (mathmath333):

for example after first time try to put x=0 then u will get 0/0 sp thats dubious, here as the deniminator had taken three steps for eliminating it ,it took 3 here

OpenStudy (anonymous):

okay,so when i continue getting0/0 i have to continue differentiation until i am sure of not getting 0

OpenStudy (anonymous):

Is that it?

OpenStudy (mathmath333):

yes most of the times

OpenStudy (anonymous):

thank you.